# Neural Audio Codecs for Prompt-Driven Universal Sound Separation

### v3

@@ 1 Introduction
We evaluate CodecSep for universal sound separation under a comprehensive set of protocols: (i) dnr-v2 for in-domain text-guided separation; (ii) five open-domain corpora—AudioCaps, ESC-50, Clotho-v2, AudioSet-eval, and VGGSound—for cross-domain generalization; (iii) three prompt granularities (fixed-stem, generic three-stem {“speech”, “music”, “sfx”}, and a universal setting with fine-grained, compositional SFX descriptions); and (iv) paraphrase robustness via unseen synonyms for speech/music. We also include an architectural ablation contrasting decoder-style source generation (CodecFormer Yip et al. (2024b)) with our Transformer masker over codec latents, plus a reconstruction diagnostic (Appendix) to probe source leakage and mixture consistency.
@@ 2 Related Work
Classical sound separation systems frequently adopt an encoder–masker–decoder design in which an encoder produces STFT-like latents, a masker predicts source-specific masks, and a decoder reconstructs waveforms. Representative models include DPTNet Chen et al. (2020b), SepFormer Subakan et al. (2021), and TDANet Li et al. (2023), the last introducing a top-down attention scheme that blends global and local attention to capture multi-scale acoustic structure. Beyond masking pipelines, several works generate waveforms directly in the time domain (Wave-UNet Stoller et al. (2018), Demucs Défossez et al. (2019, 2021)) or operate fully in the complex STFT domain with joint magnitude–phase modeling (MM-DenseLSTM Takahashi et al. (2018), DCCRN Hu et al. (2020), Spleeter Hennequin et al. (2020)), underscoring the breadth of design choices across time and frequency domains.
Moving from domain-specific separation to universal sound separation (USS), supervised systems often rely on Permutation Invariant Training (PIT) Yu et al. (2017); Kavalerov et al. (2019), while unsupervised approaches such as MixIT Wisdom et al. (2020) learn from mixtures alone. However, both typically assume a fixed maximum number of sources and emit all estimates indiscriminately, necessitating a separate identification step. Query-Guided Sound Separation (QSS) addresses this limitation by conditioning extraction on external queries—visual, exemplar audio, class labels, or text. Text queries are compact and expressive, capture high-level semantics, and avoid costly references. AudioSep Liu et al. (2024) exemplifies this direction by injecting BERT Devlin et al. (2019) or CLAP Wu et al. (2023) embeddings via FiLM Perez et al. (2018) at intermediate masker layers to steer separation toward the requested source.
@@ 3.1 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
We use DAC Kumar et al. (2023) as encoder–decoder. Following Encodec/SoundStream, DAC uses fully-convolutional encoder/decoder with the periodic Snake activation $x+\sin^{2}x$ (replacing LeakyReLU) to bias periodic audio modeling. Residual vector quantization (RVQ) compresses encoder outputs with factorized codes and $\ell_{2}$-normalized codebooks. For a 1 s audio $x(t)$ at $F_{s}{=}24$ kHz compressed to R=6000R{=}6000 bps, the encoder $Enc(\cdot)$ downsamples by M=320M{=}320 to $T{=}{F_{s}}/{M}{=}75$ frames $Z{=}[z_{t}\!\in\!\mathbb{R}^{d}]_{t=1}^{T}$ with r=R/T=80r{=}R/T{=}80 bits/frame. RVQ allocates $r_{i}{=}r/N_{q}{=}10$ bits across $N_{q}{=}8$ codebooks (size $2^{10}{=}1024$). Given $Z$, $Quant(\cdot)$ yields discrete codes $A{=}[a_{t}\!\in\![1024]^{8}]$, which map to embeddings $e_{t}{=}\sum_{i=1}^{8}e^{i}_{t}$; $Dec(\cdot)$ upsamples $E{=}[e_{t}]$ back to waveform $y(t)$.
@@
$x(t)\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
where $g(\cdot,e_{\tau})$ is a FiLM-conditioned masker that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$ given the complex STFT $X$ of audio $x(t)$ and text-embedding $e_{\tau}$.
@@
$x(t)\xrightarrow{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
@@
Operating on NAC latents $Z$ slashes dimensionality while preserving perceptual factors. For $1\,\mathrm{s}$ audio at $32\,\mathrm{kHz}$, complex STFT with N=1024N{=}1024 and hop size M=320M{=}320 samples has $T_{\mathrm{spec}}\!\approx\!100$ frames and $F=2\times 1024$ (Re+Im) scalars per frame, so $F\cdot T_{spec}\!\approx\!204{,}800$. A $16\,\mathrm{kHz}$ DAC with width d=64d{=}64 and the same $M$ yields $T\!\approx\!50$ and $d\cdot T{=}64\times 50{=}3{,}200$ ($\sim 64{\times}$ smaller), shrinking $Q/K/V$ and MLP sizes and easing self-attention. Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs. Crucially, $Enc(\cdot)$ organizes $Z$ on a discriminative, perceptually aligned manifold , making selection (masking) easier than representation learning from raw $X$. In spectrogram systems, the separator must first learn a high-level latent from $X$ (via CNN/UNet) and then separate, coupling abstraction and masking and inflating parameters/compute.
@@
Because the DAC codec induces a strong semantic prior in its latent space via residual vector quantization (RVQ) and perceptual/adversarial training, we mask the codec latents rather than generate sources from scratch as in CodecFormer. RVQ creates a coarse-to-fine hierarchy in $Z{=}Enc(x)\in\mathbb{R}^{d\times T}$: early stages capture coarse structure (e.g., low-frequency content, timbre), while later stages refine residual detail (e.g., high-frequency components, transients) Wang et al. (2023). We exploit this organization with a FiLM-conditioned transformer masker that predicts a soft mask $M_{s}\in[0,1]^{d\times T}$ and applies it element-wise, yielding source latent estimate $\tilde{Z}_{s}=M_{s}\odot Z$. In contrast to learning a generator $Sep(Z,e_{\tau}):Z\!\to\!\tilde{Z}_{s}$ as in CodecFormer, learning a mask $Mask(Z,e_{\tau}):Z\!\to\!M_{s}$ on the compact, semantically organized codec manifold both exploits the codec prior more effectively and yields a more stable optimization that converges faster. Moreover, masking in the denoised, low-dimensional codec space is fundamentally easier than masking in the high-dimensional, noisy spectrogram domain. This selection-centric design (i) constrains learning to modulation of existing latent content, (ii) avoids hallucination and reduces leakage because no new signals are synthesized, and (iii) preserves long-horizon structure (periodicity, timbre, transients) already organized by the codec, yielding stable, low-artifact separations.
@@
By design, $E\!\approx\!Z$ at the operating bitrate and $Dec(E)$ already yields high-fidelity reconstructions; because our separator is a masker (selective modulation) rather than a generator, swapping $Z\!\to\!E$ preserves the semantics needed for separation with no architectural change. While we report results on $Z$ to isolate separator performance and maintain stable optimization, we also evaluate the bitstream path by feeding reconstructed embeddings $E$ (codes-in) to the same trained masker without any fine-tuning; performance remains competitive relative to the $Z$ path. The residual gap can be narrowed with light fine-tuning the masker on $E$ or optimizing an embedding-consistency loss (cf. CodecFormer-EL) in place of, or alongside, SI-SDR: $\mathcal{L}_{\mathrm{emb}}=\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$. In deployment, the variant simply replaces the masker input with $E$ and optionally re-quantizes for codes-out.
Deployment advantages over spectrogram-domain separators.
In realistic pipelines, edge devices already run a codec and send code streams. Spectrogram models must decode $\to$ separate on $X$ $\to$ re-encode, adding latency and energy; CodecSep masks in the codec domain and can return separated code streams directly:
$\text{Edge: }Enc(\cdot)\Rightarrow\text{Code}\Rightarrow\text{Server: reconstruct }E\!\approx\!Z\ \&\ Mask(\cdot,e_{\tau})\Rightarrow\text{Code}\Rightarrow\text{Edge: }Dec(\cdot).$
Concretely, with codec costs $C_{\mathrm{enc}},C_{\mathrm{dec}}$, spectrogram separator cost $C_{\mathrm{spec}}$, and CodecSep masker cost $C_{\mathrm{mask}}$:
$\text{Code-stream input: }\ \text{AudioSep}=C_{\mathrm{dec}}{+}C_{\mathrm{spec}}{+}C_{\mathrm{enc}},\quad\text{CodecSep}=C_{\mathrm{mask}}.$
Within the separator, working on $Z/E$ (small $d,T$ with $|Z|\!\ll\!|X|$) shrinks attention/MLP activations, enabling tighter batching and lower memory bandwidth. Interface compatibility is immediate when only codes $A$ are available: perform lookup to $E$, apply the same FiLM-conditioned masker, and optionally re-quantize to emit $\hat{A}_{s}$. Overall, CodecSep (i) eliminates redundant decode/re-encode in server workflows; (ii) preserves the codec manifold (periodicity, timbre, transients) by modulating $Z/E$ instead of rebuilding $X$; and (iii) keeps conditioning non-iterative (single forward pass), yielding low-latency, high-fidelity stems at scale.
See Appendix A for a full rationale covering latent-domain masking vs. spectrogram-domain generation, the RVQ-induced semantic prior, FiLM placement inside the masker, continuous $Z$ vs. discrete $A{\to}E$ operation (and codes-in/out), and complexity/deployment comparisons.
@@ 4 Experiments
We evaluate across a controlled multi-stem corpus and multiple open-domain benchmarks. For in-domain experiments, we adapt Divide and Remaster v2 (dnr-v2) Petermann et al. (2022) from fixed-label three-stem separation to universal, prompt-driven separation by replacing source labels with natural-language queries. Speech and music are queried with broad category prompts (e.g., “speech”, “music”), while SFX stems are queried using long-form, compositional prompts (>1>1 overlapping sources) synthesized from FSD50K’s hierarchical annotations Fonseca et al. (2022), combining fine-grained classes with parent categories (e.g., “dog barking, Animal, engine rumbling, motor vehicle”). To assess cross-domain generalization, we form three-source mixtures on AudioCaps Kim et al. (2019) (used for both training and testing in our open-domain setting) and construct test-only three-source mixtures from ESC-50 Piczak , Clotho-v2 Drossos et al. (2020), AudioSet-eval Gemmeke et al. (2017), and VGGSound Chen et al. (2020a). Dataset construction details, clip durations, split statistics, and segmentation rules are provided in the Appendix B.
@@
We compare CodecSep against representative spectrogram- and codec-domain baselines—TDANet Li et al. (2023); Pons et al. (2024), CodecFormer Yip et al. (2024b), SDCodec Bie et al. (2024), and the text-guided spectrogram model AudioSep Liu et al. (2024). We report objective signal fidelity via scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019) and perceptual quality via ViSQOL Chinen et al. (2020), which measures spectro–temporal similarity between the estimate $\tilde{x}$ and reference $x$ and maps it to a 1–5 MOS-LQO scale. Following prior work (e.g., SDCodec Bie et al. (2024)), we use ViSQOL as a proxy for perceptual listening quality and complement it with a human MOS-LQS study comparing real-world outputs from CodecSep (trained on dnr-v2) and the publicly released AudioSep. To quantify efficiency in a hardware-agnostic manner, we report multiply–accumulate operations (MACs) using torchprofile [^1] https://github.com/zhijian-liu/torchprofile under matched input durations (2 s) and batching (batch size 2).
@@
Unless otherwise stated, the DAC codec Kumar et al. (2023) and CLAP text encoder Wu et al. (2023) remain frozen; we train the FiLM-conditioned transformer masker and the lightweight query network end-to-end with an Adam optimizer Kingma and Ba (2017) and a plateau-based learning-rate schedule Mukherjee et al. (2019). We produce two variants of CodecSep, trained separately on dnr-v2 and on AudioCaps, to study distributional effects; we denote them with the suffixes “+dnr-v2” and “+AudioCaps”. For fair comparison, TDANet and CodecFormer are re-trained from scratch on dnr-v2 using our setup, AudioSep is evaluated both as a public checkpoint and when re-trained under matched protocols, and SDCodec is used as released by the authors. To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through a full-band stereo-capable $48$ kHz EnCodec during both training and inference. Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix C.
@@
AudioSep SI-SDR $2.46^{\pm 4.06}$ $4.92^{\pm 4.21}$ $0.34^{\pm 5.39}$
(zero-shot) ViSQOL $2.86^{\pm 0.63}$ $3.11^{\pm 0.56}$ $\mathbf{2.63^{\pm 0.77}}$
AudioSep + dnr-v2 SI-SDR $5.55^{\pm 2.89}$ $7.68^{\pm 3.0}$ $4.66^{\pm 3.68}$
ViSQOL $2.59^{\pm 0.57}$ $2.49^{\pm 0.37}$ $2.32^{\pm 0.7}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.15^{\pm 3.29}}$ $\mathbf{9.97^{\pm 2.92}}$ $\mathbf{0.89^{\pm 4.22}}$
ViSQOL $\mathbf{2.86^{\pm 0.57}}$ $\mathbf{3.14^{\pm 0.45}}$ $2.33^{\pm 0.73}$
CodecSep + dnr-v2 SI-SDR $0.16^{\pm 3.55}$ $8.26^{\pm 2.60}$ $0.96^{\pm 4.20}$
(codes in : codes out, zero-shot) ViSQOL $2.52^{\pm 0.52}$ $2.97^{\pm 0.44}$ $2.33^{\pm 0.67}$
CodecSep + dnr-v2 SI-SDR $6.72^{\pm 2.77}$ $1.95^{\pm 2.84}$ $6.75^{\pm 3.83}$
(ablate Masker) ViSQOL $2.48^{\pm 0.58}$ $2.58^{\pm 0.50}$ $2.08^{\pm 0.74}$

### v4

@@ 2 Related Work
Classical sound separation systems frequently adopt an encoder–masker–decoder design in which an encoder produces STFT-like latents, a masker predicts source-specific masks, and a decoder reconstructs waveforms. Representative models include DPTNet Chen et al. (2020b), SepFormer Subakan et al. (2021), and TDANet Li et al. (2023), the last introducing a top-down attention scheme that blends global and local attention to capture multi-scale acoustic structure. Beyond masking pipelines, several works generate waveforms directly in the time domain (Wave-UNet Stoller et al. (2018), Demucs Défossez et al. (2019, 2021)) or operate fully in the complex STFT domain with joint magnitude–phase modeling (MM-DenseLSTM Takahashi et al. (2018), DCCRN Hu et al. (2020), Spleeter Hennequin et al. (2020)), underscoring the breadth of design choices.
Moving from domain-specific separation to universal sound separation (USS), supervised systems typically rely on Permutation Invariant Training (PIT) Yu et al. (2017); Kavalerov et al. (2019), while unsupervised methods such as MixIT Wisdom et al. (2020) learn directly from mixtures. Both paradigms assume a fixed maximum number of sources and output all estimates indiscriminately, requiring a post-hoc identification step (cf. Appendix A for detailed failure modes). A recent PIT-trained USS model is Sudo rm-rf Tzinis et al. (2022a), a parameter-efficient time-domain separator that is based on ConvTasNet-style encoder–masker–decoder architecture with adaptive encoder/decoder modules. It downsamples input waveform to STFT-like latents before separation and is PIT-trained to separate mixtures with up to four sources. Query-Guided Sound Separation (QSS) addresses this limitation of PIT or MixIT models by conditioning extraction on external queries—visual cues, audio, class labels, or text. Text queries are compact, expressive, and capture high-level semantics without requiring additional reference signals. AudioSep Liu et al. (2024) follows this direction by injecting BERT Devlin et al. (2019) or CLAP Wu et al. (2023) embeddings via FiLM Perez et al. (2018) at intermediate masker layers to steer separation toward the query source. BiModalSS Mahmud et al. (2024) extends AudioSep with attention-based conditioning and more efficient training strategies. FiLM-conditioned variants of Sudo rm-rf Tzinis et al. (2022b, 2023) have also been explored for class-guided separation using one-hot or multi-hot conditioning vectors; however, such label-based conditioning does not extend naturally to open-domain text queries and cannot handle unseen classes.
@@ 3.1 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
We use DAC Kumar et al. (2023) as encoder–decoder. Following Encodec/SoundStream, DAC uses fully-convolutional encoder/decoder with the periodic Snake activation $x+\sin^{2}x$ (replacing LeakyReLU) to bias periodic audio modeling. Residual vector quantization (RVQ) compresses encoder outputs with factorized codes and $\ell_{2}$-normalized codebooks. For a 1 s audio $x(t)$ at $F_{s}{=}24$ kHz compressed to R=6000R{=}6000 bps, the encoder $Enc(\cdot)$ downsamples by M=320M{=}320 to $T{=}{F_{s}}/{M}{=}75$ frames of latents $Z{=}[z_{t}\!\in\!\mathbb{R}^{d}]_{t=1}^{T}$ ($d$: channel width) with r=R/T=80r{=}R/T{=}80 bits/frame. RVQ allocates $r_{i}{=}r/N_{q}{=}10$ bits across $N_{q}{=}8$ codebooks (size $2^{10}{=}1024$). Given $Z$, $Quant(\cdot)$ yields discrete codes $A{=}[a_{t}\!\in\![1024]^{8}]$, which map to embeddings $e_{t}{=}\sum_{i=1}^{8}e^{i}_{t}$; $Dec(\cdot)$ upsamples $E{=}[e_{t}]$ back to waveform $y(t)$.
@@
$x(t)\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,Spec(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
where $Spec(\cdot,e_{\tau})$ is a FiLM-conditioned masker that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ ($F$: frequency bins, $T_{spec}$: spectrogram frames) and a phase residual $\angle\hat{M}_{s}$ given the complex STFT $X$ of audio $x(t)$ and text-embedding $e_{\tau}$.
@@
$x(t)\xrightarrow[\text{DAC}]{\,Enc(\cdot)\,}Z\!\in\!\mathbb{R}^{d\times T}\;\xrightarrow{\,Mask(Z,e_{\tau})\,}\;\tilde{Z}_{s}=M_{s}\!\odot\!Z\;\xrightarrow[\text{DAC}]{\,Dec(\cdot)\,}\;\tilde{y}_{s}(t),$
@@
Operating on NAC latents $Z$ slashes dimensionality while preserving perceptual factors. For $1\,\mathrm{s}$ audio at $32\,\mathrm{kHz}$, complex STFT with N=1024N{=}1024 and hop size M=320M{=}320 samples has $T_{\mathrm{spec}}\!\approx\!100$ frames and $F=2\times 1024$ (Re+Im) scalars per frame, so $F\cdot T_{spec}\!\approx\!204{,}800$. A $16\,\mathrm{kHz}$ DAC with width d=64d{=}64 and the same $M$ yields $T\!\approx\!50$ and $d\cdot T{=}64\times 50{=}3{,}200$ ($\sim 64{\times}$ smaller), shrinking $Q/K/V$ and MLP sizes and easing self-attention. Similarly, for $32\,\mathrm{kHz}$ NACs like EnCodec, $T\!\approx\!50$ with d=128d{=}128, so attention/MLPs still operate on $\sim 32{\times}$ smaller latents than complex STFTs. Crucially, $Enc(\cdot)$ organizes $Z$ on a discriminative, perceptually aligned manifold , making selection (masking) easier than representation learning from raw $X$. In spectrogram systems $Spec(.)$, the separator must first learn a high-level latent from $X$ (via CNN/UNet) and then separate, coupling abstraction and masking and inflating parameters/compute. Waveform separators $Wave(.)$ such as Sudo rm-rf Tzinis et al. (2022a, b, 2023) likewise downsample the waveform into STFT-like intermediate latents via 1D convolutions before encoding, resulting in latents with similar dimensionality and thus inheriting the same challenges as spectrogram systems.
@@
Because the DAC codec induces a strong semantic prior in its latent space via residual vector quantization (RVQ) and perceptual/adversarial training, we mask the codec latents rather than generate sources from scratch as in CodecFormer. RVQ creates a coarse-to-fine hierarchy in $Z{=}Enc(x)\in\mathbb{R}^{d\times T}$: early stages capture coarse structure (e.g., low-frequency content, timbre), while later stages refine residual detail (e.g., high-frequency components, transients) Wang et al. (2023). We exploit this organization with a FiLM-conditioned transformer masker that predicts a soft mask $M_{s}\in[0,1]^{d\times T}$ and applies it element-wise, yielding source latent estimate $\tilde{Z}_{s}=M_{s}\odot Z$. In contrast to learning a generator $Gen(Z,e_{\tau}):Z\!\to\!\tilde{Z}_{s}$ as in CodecFormer, learning a mask $Mask(Z,e_{\tau}):Z\!\to\!M_{s}$ on the compact, semantically organized codec manifold both exploits the codec prior more effectively and yields a more stable optimization that converges faster. Moreover, masking in the denoised, low-dimensional codec space is fundamentally easier than masking in the high-dimensional, noisy spectrogram domain. This selection-centric design (i) constrains learning to modulation of existing latent content, (ii) avoids hallucination and reduces leakage because no new signals are synthesized, and (iii) preserves long-horizon structure (periodicity, timbre, transients) already organized by the codec, yielding stable, low-artifact separations.
@@
By design, $E\!\approx\!Z$ at the operating bitrate and $Dec(E)$ already yields high-fidelity reconstructions; because our separator is a masker (selective modulation) rather than a generator, swapping $Z\!\to\!E$ preserves the semantics needed for separation with no architectural change. While we report results on $Z$ to isolate separator performance and maintain stable optimization, we also evaluate the bitstream path by feeding reconstructed embeddings $E$ (codes-in) to the same trained masker without any fine-tuning; performance remains competitive relative to the $Z$ path. The residual gap can be narrowed with light fine-tuning the masker on $E$ or optimizing an embedding-consistency loss (cf. CodecFormer-EL) in place of, or alongside, SI-SDR: $\mathcal{L}_{\mathrm{emb}}=\sum_{s}\|\tilde{E}_{s}-Z_{s}\|_{1}$. In deployment, the variant simply replaces the masker input with $E$ and optionally re-quantizes for codes-out as,
$x(t)\;\xrightarrow[\text{On Edge}]{\,Quant(Enc(\cdot))\,}\;A\;\xrightarrow[\text{Codes In}]{\,lookup(A)\,}\;E\approx Z\;\xrightarrow[\text{On Server}]{\,Mask(E,e_{\tau})\,}\;\tilde{E}_{s}=M_{s}\!\odot\!E\;\xrightarrow[\text{Codes Out}]{\,Quant(\tilde{E}_{s})\,}\;\hat{A}_{s}.$
Deployment advantages over audio-stream separators.
In realistic pipelines, edge devices already run a codec and transmit code streams rather than raw audio. Traditional spectrogram-based $Spec(.)$ and waveform-based $Wave(.)$ separators, however, operate on the audio stream: they first convert audio to STFT or STFT-like representations (often via 1D convolutions), and then must decode $\to$ separate on $X$ $\to$ re-encode, incurring additional latency and energy cost. In contrast, CodecSep performs masking directly in the codec domain and can output code streams without any decode–re-encode cycle. Concretely, with codec costs $C_{{Enc}},C_{{Dec}}$, spectrogram or audio-stream separator (AudioSep) cost $C_{{Spec}}$, and CodecSep masker cost $C_{{Mask}}$:
$\text{Compute Cost for Code-stream input: }\ \text{AudioSep}=C_{{Dec}}{+}C_{{Spec}}{+}C_{{Enc}},\quad\text{CodecSep}=C_{{Mask}}.$
We treat the codebook lookup $C_{{lookup}}$ and quantization $C_{{Quant}}$ costs as negligible ($\approx 0$) and omit the CLAP text-encoder cost since it is shared across all models. Figure 2 illustrates a typical edge–server deployment and compares compute requirements for conventional audio-stream separators (audio in $\to$ codes out) versus CodecSep’s code-stream separator (codes in $\to$ codes out). As shown, the CodecSep masker operates on $Z/E$ with small $(d,T)$ where $|Z|\ll|X|$, dramatically reducing attention and MLP activations and enabling tighter batching and lower memory bandwidth. Interface compatibility is immediate when only codes $A$ are available: perform a lookup to obtain $E$, apply the FiLM-conditioned masker, and optionally re-quantize to produce $\hat{A}_{s}$. CodecSep thus eliminates redundant decode/re-encode cycles in server workflows yielding low-latency, high-fidelity separation at scale. See Appendix B for a full discussion covering all of the aforementioned rationale in §3.1.
Figure 2: Typical edge–server deployment comparing compute requirements of conventional audio-stream separators (audio in $\to$ codes out) versus CodecSep discrete inference (codes in $\to$ codes out).
@@ 4 Experiments
We evaluate across a controlled multi-stem corpus and multiple open-domain benchmarks. For in-domain experiments, we adapt Divide and Remaster v2 (dnr-v2) Petermann et al. (2022) from fixed-label three-stem separation to universal, prompt-driven separation by replacing source labels with natural-language queries. Speech and music are queried with broad category prompts (e.g., “speech”, “music”), while SFX stems are queried using long-form, compositional prompts ($\geq 2$ overlapping sources) synthesized from FSD50K’s hierarchical annotations Fonseca et al. (2022), combining fine-grained classes with parent categories (e.g., “dog barking, Animal, engine rumbling, motor vehicle”). To assess cross-domain generalization, we form three-source mixtures on AudioCaps Kim et al. (2019) (used for both training and testing in our open-domain setting) and construct test-only three-source mixtures from ESC-50 Piczak , Clotho-v2 Drossos et al. (2020), AudioSet-eval Gemmeke et al. (2017), and VGGSound Chen et al. (2020a). Dataset construction details, clip durations, split statistics, and segmentation rules are provided in the Appendix C.
@@
We compare CodecSep against representative spectrogram-, waveform- and codec-domain baselines—TDANet Li et al. (2023); Pons et al. (2024), Sudo rm-rf Tzinis et al. (2022a), CodecFormer Yip et al. (2024b), SDCodec Bie et al. (2024), and the text-guided audio stream separators AudioSep Liu et al. (2024), BiModalSS Mahmud et al. (2024) and Sudo rm-rf + FiLM Tzinis et al. (2022b, 2023). We report objective signal fidelity via scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019) and perceptual quality via ViSQOL Chinen et al. (2020), which measures spectro–temporal similarity between the estimate $\tilde{x}$ and reference $x$ and maps it to a 1–5 MOS-LQO scale. Following prior work (e.g., SDCodec Bie et al. (2024)), we use ViSQOL as a proxy MOS score for perceptual listening quality and complement it with a human MOS-LQS study comparing real-world outputs from CodecSep (trained on dnr-v2) and the publicly released AudioSep. To quantify efficiency, we report multiply–accumulate operations (MACs), inference time, memory footprint using torchinfo [^1] https://github.com/tyleryep/torchinfo under matched input durations (2 s) and batching (batch size 2) on a 2424GB NVIDIA A30 GPU. Details on evaluation workflow for each benchmark are deferred to Appendix D.
@@
Unless otherwise stated, the DAC codec Kumar et al. (2023) and CLAP text encoder Wu et al. (2023) remain frozen; we train the FiLM-conditioned transformer masker and the lightweight query network end-to-end with an Adam optimizer Kingma and Ba (2017) and a plateau-based learning-rate schedule Mukherjee et al. (2019). We produce two variants of CodecSep, trained separately on dnr-v2 and on AudioCaps, to study distributional effects; we denote them with the suffixes “+dnr-v2” and “+AudioCaps”. For fair comparison, the 3-stem versions of TDANet, Sudo rm-rf, and CodecFormer are re-trained from scratch on dnr-v2 using our setup. AudioSep is evaluated both as the publicly released checkpoint and when re-trained under matched protocols. We similarly re-train Sudo rm-rf+FiLM under the same matched settings. Pretrained checkpoints of BiModalSS and SDCodec are used as released by the authors. To reflect realistic deployments where signals traverse compression pipelines, inputs to non-codec baselines are passed through a full-band stereo-capable $48$ kHz EnCodec during both training and inference. Full hyper-parameters, iteration schedules, batch configurations, and hardware details are deferred to the Appendix E.
@@
AudioSep SI-SDR $2.5^{\pm 4.06}$ $4.9^{\pm 4.21}$ $0.3^{\pm 5.39}$
(zero-shot) ViSQOL $2.9^{\pm 0.63}$ $3.1^{\pm 0.56}$ $\mathbf{2.6^{\pm 0.77}}$
Bimodal SI-SDR $6.8^{\pm 2.73}$ $1.8^{\pm 2.78}$ $6.36^{\pm 3.57}$
(zero-shot) ViSQOL $2.5^{\pm 0.57}$ $2.6^{\pm 0.51}$ $2.3^{\pm 0.72}$
AudioSep + dnr-v2 SI-SDR $5.6^{\pm 2.89}$ $7.7^{\pm 3.0}$ $4.7^{\pm 3.68}$
ViSQOL $2.6^{\pm 0.57}$ $2.5^{\pm 0.37}$ $2.3^{\pm 0.7}$
Sudo rm-rf + FiLM + dnr-v2 SI-SDR $6.7^{\pm 2.62}$ $2.0^{\pm 2.76}$ $6.6^{\pm 3.71}$
ViSQOL $2.7^{\pm 0.59}$ $2.9^{\pm 0.45}$ $2.3^{\pm 0.72}$
CodecSep + dnr-v2 SI-SDR $\mathbf{1.2^{\pm 3.29}}$ $\mathbf{10.0^{\pm 2.92}}$ $\mathbf{0.9^{\pm 4.22}}$
ViSQOL $\mathbf{2.9^{\pm 0.57}}$ $\mathbf{3.2^{\pm 0.45}}$ $2.3^{\pm 0.73}$
CodecSep + dnr-v2 SI-SDR $0.2^{\pm 3.55}$ $8.3^{\pm 2.60}$ $1.0^{\pm 4.20}$
(codes in : codes out, zero-shot) ViSQOL $2.5^{\pm 0.52}$ $3.0^{\pm 0.44}$ $2.3^{\pm 0.67}$
CodecSep + dnr-v2 SI-SDR $6.8^{\pm 2.77}$ $2.0^{\pm 2.84}$ $6.8^{\pm 3.83}$
(ablate Masker) ViSQOL $2.5^{\pm 0.58}$ $2.6^{\pm 0.50}$ $2.1^{\pm 0.74}$